# On the gcd graphs over polynomial rings and related topics

### v1

@@ Abstract.
Gcd-graphs over the ring of integers modulo $n$ are a natural generalization of unitary Cayley graphs. The study of these graphs has foundations in various mathematical fields, including number theory, ring theory, and representation theory. Using the theory of Ramanujan sums, it is known that these gcd-graphs have integral spectra; i.e., all their eigenvalues are integers. In this work, inspired by the analogy between number fields and function fields, we define and study gcd-graphs over polynomial rings with coefficients in finite fields. We establish some fundamental properties of these graphs, emphasizing their analogy to their counterparts over $\operatorname{\mathbb{Z}}.$
@@
Cayley graphs, Function fields, Ramanujan sums, Symmetric algebras, Integral graphs, Prime graphs, Perfect graphs.
@@ 1. Introduction
Let $n$ be a positive integer. The unitary Cayley graph on the ring of integers modulo $n$ is defined as the Cayley graph $G_{n}={\rm Cay}(\operatorname{\mathbb{Z}}/n,U_{n})$ where $\operatorname{\mathbb{Z}}/n$ is the ring of integers modulo $n$ and $U_{n}=(\operatorname{\mathbb{Z}}/n)^{\times}$ its unit group. More specifically, $G_{n}$ is equipped with the following data.
@@
    The vertex set of $G_{n}$ is $\operatorname{\mathbb{Z}}/n$.
@@
The unitary Cayley graph $G_{n}$ was first formally introduced in [18] even though we can trace it back to the work of Evans and Erdős [11]. Due to its elegance and simplicity, the unitary Cayley graph has been further studied and generalized by various works in the literature. For example, [1] studies the unitary Cayley graph of a finite commutative ring. [15] generalizes this study further to finite rings which are not necessarily commutative. In [26], a subset of the authors studies various arithmetic and graph-theoretic properties of the $p$-unitary Cayley graph as defined by [27, 28] (when $p=1$, these graphs are precisely unitary Cayley graphs). We refer the interested readers to [3, 8, 18, 26] and the references therein for some further topics in this line of research.
As explained in [18, Section 4], a particularly intriguing arithmetical property of $G_{n}$ is that its spectrum can be described by the theory of circulant matrices and Ramanujan sums. A consequence of this fact is that the unitary Cayley graph has an integral spectrum; i.e., all of its eigenvalues are integers. In [18], the authors note that the unitary Cayley graph is not the sole graph exhibiting an integral spectrum. They identify a closely related family of graphs sharing this characteristic, known as gcd-graphs which we now recall. Let $D=\{d_{1},d_{2},\ldots,d_{k}\}$ be a set of proper divisors of $n$. The gcd-graph $G_{n}(D)$ is the graph with the following data.
@@
    The vertex set of $G_{n}(D)$ is $\operatorname{\mathbb{Z}}/n.$
@@
We remark that the unitary Cayley graph $G_{n}$ is nothing but $G_{n}(\{1\}).$ It is known that the spectrum of $G_{n}(D)$ is a summation of various Ramanujan sums (see [18, Section 4]). In particular, all of its eigenvalues are integers. It turns out that, the converse is also true.
Theorem 1.1.
(See [29, Theorem 7.1]) A $\operatorname{\mathbb{Z}}/n$-circulant graph $G$ is an integral graph if and only if $G=G_{n}(D)$ for some $D.$
Let $\operatorname{\mathbb{F}}_{q}$ be a finite field. As observed by Andre Weil in [30], there is a strong analogy between the ring $\operatorname{\mathbb{Z}}$ of integers and the ring $\operatorname{\mathbb{F}}_{q}[x]$ of polynomials with coefficients in $\operatorname{\mathbb{F}}_{q}$ (and more generally, between number fields and function fields). Consequently, it is of reasonable interest to define and investigate gcd-graphs over $\mathbb{F}_{q}[x]$. Fortunately, the analogous definition of gcd-graphs over $\mathbb{F}_{q}[x]$ is relatively straightforward. More specifically, let $f\in\operatorname{\mathbb{F}}_{q}[x]$ be a non-zero element in $\operatorname{\mathbb{F}}_{q}[x]$. Let $D=\{f_{1},f_{2},\ldots,f_{k}\}$ be a subset of the set of divisors $\operatorname{{\rm Div}}(f)$ of $f$; i.e, $f_{i}\mid f$ for all $1\leq i\leq k.$ Let
$S_{D}=\{g\in\operatorname{\mathbb{F}}_{q}[x]/f|\gcd(g,f)\in D\}.$
@@
We denote by $G_{f}(D)$ the gcd-Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S_{D}).$ More precisely, it is the graph equipped with the following data
@@
    The vertex set of $G_{f}(D)$ is the finite ring $\operatorname{\mathbb{F}}_{q}[x]/f.$
@@
Remark 1.3.
In general, the greatest common divisor is only defined up to associates. In our case, however, we can make a canonical choice for the great common divisor by requiring it to be a monic polynomial. Therefore, unless we explicitly state, we will assume throughout this article that all involved polynomials are monic.
@@
Fig. 1 shows the graph $G_{f}(D)$ where $f=x(x+1)\in\operatorname{\mathbb{F}}_{3}[x]$ and $D=\{x,x+1\}$. It is a regular graph of degree $4$
Figure 1. The gcd-Cayley graph $G_{x(x+1)}(\{x,x+1\})$
1.1. Outline
In this article, we study some foundational properties of gcd-graphs defined over $\mathbb{F}_{q}[x]$. In Section 2, we recall some standard definitions in graph theory that will be used throughout the article. Section 3 investigates the question of when a gcd-graph is connected and anti-connected. In Section 4, we provide the necessary and sufficient conditions for a gcd-graph to be bipartite. Section 5 explores another graph-theoretic property of gcd-graphs: their prime-property. Although we are not able to provide a complete answer, we present some rather general conditions under which this property holds. Section 6 examines the spectra of gcd-graphs using symmetric algebras and Ramanujan sums. We find it quite remarkable that the explicit formulas for these spectra are identical to those that appeared in the number field case. In Section 7, we provide some sufficient conditions for gcd-graphs to be perfect. Finally, in Section 8, we investigate whether a given graph can be realized as an induced subgraph of a gcd-graph. As a by-product, we prove an analogous result to a theorem of Erdős and Evans in the number field case.
1.2. Code
@@ 2. Background from graph theory
$K_{n}$ is the graph on $n$ vertices which are pairwisely adjacent.
@@
Definition 2.3 (Wreath product).
Let $G,H$ be two graphs. We define the wreath product of $G$ and $H$ as the graph $G*H$ with the following data
@@
    The vertex set of $G*H$ is the Cartesian product $V(G)\times V(H)$,
@@
    $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent in $G*H$ if either $(g,g^{\prime})\in E(G)$ or $g=g^{\prime}$ and $(h,h^{\prime})\in E(H)$.
Definition 2.4 (Graph morphism).
Let $G_{1}$ and $G_{2}$ be two graphs. We define a graph morphism between $G_{1}$ and $G_{2}$ to be a map from $V(G_{1})$ to $V(G_{2})$ $f$ that preserves edges. More precisely, if $u,v\in V(G_{1})$ are adjacent in $G_{1}$, then $f(u),f(v)$ are adjacent in $G_{2}$.
Definition 2.5 (Induced subgraph).
Let $G_{1}$ and $G_{2}$ be two graphs. We say that $G_{1}$ is an induced subgraph of $G_{2}$ if there exists a graph morphism $f:G_{1}\to G_{2}$ such that the following conditions hold.
@@
    For $u,v\in G_{1}$, $u,v$ are adjacent if and only if $f(u)$ and $f(v)$ are adjacent in $G_{2}.$
3. Connectedness of gcd graphs
In this section, we will give the necessary and sufficient conditions for $G_{f}(D)$ to be connected. Unlike the number field case where these conditions are rather straightforward, the function field case is a bit more complicated. This is partially due to the fact that the unitary Cayley graph on $\operatorname{\mathbb{F}}_{q}[x]/f$ is not always connected. We start our discussion with the following simple observation.
@@ 3. Connectedness of gcd graphs
Let $f_{0}=\gcd(f_{1},\ldots,f_{k})$. We observe that if two vertices $u$ and $v$ are adjacent then $f_{0}$ divides $u-v$. Now since $G_{f}(D)$ is connected, there is a path connecting 00 and $1$. This implies that $f_{0}$ divides $1-0=1$. Hence $\gcd(f_{1},\ldots,f_{k})=1$. ∎
In the case of $\operatorname{\mathbb{Z}}$, the converse of 3.1 is true as well; i.e. the gcd-graph $G_{n}(D)$, where $D=\{d_{1},d_{2},\ldots,d_{k}\}$ is connected if and only if ${\rm gcd}(d_{1},d_{2},\ldots,d_{k})=1.$ In the case of $\operatorname{\mathbb{F}}_{q}[x]$, this condition is not sufficient. For example, let $D=\{1\}$, $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2},f=x(x+1).$ Then $\operatorname{\mathbb{F}}_{q}[x]/f\cong\operatorname{\mathbb{F}}_{2}\times\operatorname{\mathbb{F}}_{2}.$ In this case
@@
We need to show that $R=\langle S_{D}\rangle$ the abelian group generated by $S_{D}$. Let $a\in R=\operatorname{\mathbb{F}}_{q}[x]/f.$ Because $\gcd(f_{1},f_{2},\ldots,f_{k})=1$, we can find $a_{1},a_{2},\ldots,a_{k}\in R$ such that
@@
Since $G_{f}(\{1\})$ is connected, for each $1\leq i\leq d$, we can find write
@@
By definition, $\gcd(s_{ij}f_{i},f)=f_{i}\in D$. This shows that $a\in\langle S_{D}\rangle$. Since this is true for all $a$, we conclude that $R=\langle S_{D}\rangle$. ∎
@@
Proposition 3.3.
(See [8, Lemma 4.33]) $G_{f}(\{1\})$ is not connected if and only if $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$
@@
Let $g$ be a divisor of $f$ and $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/g$ be the canonical projection map. Let
$\overline{D}=\{\gcd(f_{i},g)|1\leq i\leq k\}.$
@@
Let $\bar{a},\bar{b}$ be two vertices in $G_{g}(\overline{D}).$ Since $\Phi_{f,g}$ is surjective, we can find $a,b\in\operatorname{\mathbb{F}}_{q}[x]/f$ such that $\Phi_{f,g}(a)=\bar{a},\Phi_{f,g}(b)=\bar{b}.$ Since $G_{f}(D)$ is connected, there is a path $P$ from $a$ to $b.$ By 3.6, $\Phi_{f,g}(P)$ is a path from $\bar{a}$ to $\bar{b}.$ This shows that $G_{g}(\overline{D})$ is connected as well. ∎
@@
Suppose that $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$ Let $\Phi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/(x(x+1))$ be the canonical projection map and $\bar{D}$ as described above. Then $G_{f}(D)$ is connected if and only if
@@
    the graph $G_{x(x+1)}(\bar{D})$ is connected. This condition is equivalent to $|\overline{D}\setminus\{x(x+1)\}|\geq 2.$
@@
By 3.7, we know that if $G_{f}(D)$ is connected then $(1)$ and $(2)$ holds. Conversley, let us assume that $(1)$ and $(2)$ both hold. We will show that $G_{f}(D)$ is connected. This is equivalent to showing that $R:=\operatorname{\mathbb{F}}_{q}[x]/f=\langle S_{D}\rangle.$
@@
We first claim that if $g\in R$ such that $x(x+1)\mid g$ then $g=s_{1}+s_{2}$ where $s_{1},s_{2}\in R^{\times}.$ For an element $s\in R$, $s$ is a unit if and only $\Phi^{\text{ss}}(s)\in(R^{\text{ss}})^{\times}$ where $R^{\text{ss}}=R/J(R)$ is the semisimplification of $R$ (see [8, Proposition 4.30]). Therefore, for this statement, we can assume that $R=R^{\text{ss}}$; namely $f$ is a squarefree polynomial. If we write $f=x(x+1)f_{1}$ where $\gcd(x(x+1),f_{1})=1$ then we have the isomorphism
$R\cong\operatorname{\mathbb{F}}_{2}[x]/(x(x+1))\times\operatorname{\mathbb{F}}_{2}[x]/f_{1}.$
Under this isomorphism, $g$ is sent to $(0,g_{1})$ where $g_{1}\in\operatorname{\mathbb{F}}_{2}[x]/f_{1}.$ Since $F_{2}[x]/f_{1}$ is a product of fields of order bigger than $2$, every element in it can be written as the sum of two units; say $g_{1}=t_{1}+t_{2}$ where $t_{1},t_{2}\in(\operatorname{\mathbb{F}}_{2}[x]/f)^{\times}.$ We then see that $g=s_{1}+s_{2}$ where $s_{1}=(1,t_{1})$ and $s_{2}=(1,t_{2})$. By definition $s_{1},s_{2}\in R^{\times}.$
@@
Since $s_{1},s_{2}\in R^{\times}$, $s_{1i}f_{i},s_{2i}f_{i}\in S_{D}.$ This shows that $g\in\langle S_{D}\rangle.$
Finally, let $g$ now be an arbitrary element in $R$. We claim that $g\in\langle S_{D}\rangle.$ By our assumption, the graph $G_{x(x+1)}(S_{\bar{D}})$ is connected, 00 and $\Phi_{f,x(x+1)}(g)$ are connected by a path. Consequently, we can write
@@
The proof of 3.8 also implies that for the case $\operatorname{\mathbb{F}}_{2}[x]$ and $x(x+1)\mid f$, the map $\Phi_{f,g}:\operatorname{\mathbb{F}}_{2}[x]/f\to\operatorname{\mathbb{F}}_{2}[x]/(x(x+1))$ has the property that $a$ and $b$ belong the same connected component in $G_{f}(D)$ if and only $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ belong to the same connected component in $G_{x(x+1)}(\overline{D}).$ In fact, by 3.7, if $a$ and $b$ are connected by a path then $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ are connected by a path. Conversely, if $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ are connected by a path in $G_{x(x+1)}(\overline{D})$, we can write
@@
where $n_{i}\in\operatorname{\mathbb{Z}}$ and $\gcd(h_{1},f)\in D$. Because $x(x+1)R\subseteq\langle S_{D}\rangle$, we conclude that that $a-b\in\langle S_{D}\rangle.$ By definition, $a$ and $b$ are connected by a path in $G_{f}(D)$.
We conclude that $G_{f}(D)$ and $G_{x(x+1)}(\overline{D})$ has the same number of connected component. In particular, the number of connected component in $G_{f}(D)$ is at most $2$

### v2

# On the gcd-graphs over polynomial rings

@@ Abstract.
[…]
@@
Gcd-graphs, Cayley graphs, prime graphs, perfect graphs, connected graphs, bipartite graphs, Ramanujan sums, integral spectra
@@ 1. Introduction
Let $n$ be a positive integer. The unitary Cayley graph on the ring of integers modulo $n$ is defined as the Cayley graph $G_{n}={\rm Cay}(\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}},U_{n})$ where $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$ is the ring of integers modulo $n$ and $U_{n}=(\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}})^{\times}$ its unit group. More specifically, $G_{n}$ is equipped with the following data:
@@
    The vertex set of $G_{n}$ is $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$;
@@
The unitary Cayley graph $G_{n}$ was first formally introduced in [19] even though we can trace it back to the work of Evans and Erdős [12]. Due to its elegance and simplicity, the unitary Cayley graph has been further studied and generalized by various works in the literature. For example, [1] studies the unitary Cayley graph of a finite commutative ring. In [16], the authors generalize this study further to finite rings which are not necessarily commutative. In [27], a subset of the authors studies various arithmetic and graph-theoretic properties of a subclass of $k$-unitary Cayley graphs as defined in [29, 31] (when $k=1$, these graphs are precisely unitary Cayley graphs). We refer the interested readers to [4, 9, 19, 27] and the references therein for some further topics in this line of research.
As explained in [19, Section 4], a particularly intriguing arithmetical property of $G_{n}$ is that its spectrum can be described by the theory of circulant matrices and Ramanujan sums. A consequence of this fact is that the unitary Cayley graph has an integral spectrum; i.e., all of its eigenvalues are integers. In [19], the authors note that the unitary Cayley graph is not the sole graph exhibiting an integral spectrum. They identify a closely related family of graphs sharing this characteristic, known as gcd-graphs which we now recall. Let $D=\{d_{1},d_{2},\ldots,d_{k}\}$ be a set of proper divisors of $n$. The gcd-graph $G_{n}(D)$ is the graph with the following data.
@@
    The vertex set of $G_{n}(D)$ is $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}.$
@@
We remark that the unitary Cayley graph $G_{n}$ is nothing but $G_{n}(\{1\}).$ We also note that the gcd-graph $G_{n}(D)$ is loopless since $D$ does not contain $n$. Additionally, the graph $G_{n}(D)$ is undirected since $\gcd(a-b,n)=\gcd(b-a,n)$. It is known that the spectrum of $G_{n}(D)$ is a summation of various Ramanujan sums (see [19, Section 4]). In particular, all of its eigenvalues are integers. It turns out that, the converse is also true.
Theorem 1.1 (See [32, Theorem 7.1]).
An $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$-circulant graph $G$ is an integral graph if and only if $G=G_{n}(D)$ for some $D.$
[…]
$S_{D}=\{g\in\operatorname{\mathbb{F}}_{q}[x]/(f):\gcd(g,f)\in D\}.$
Remark 1.2.
In general, the greatest common divisor is only defined up to associates. In our case, however, we can make a canonical choice for the great common divisor by requiring it to be a monic polynomial. Therefore, unless we explicitly state the contrary, we will assume throughout this article that all involved polynomials are monic.
@@
The gcd-graph $G_{f}(D)$ is the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/(f),S_{D}).$ More precisely, it is the graph equipped with the following data:
@@
    The vertex set of $G_{f}(D)$ is the finite ring $\operatorname{\mathbb{F}}_{q}[x]/(f)$,
@@
[…]
@@
Fig. 1 shows the graph $G_{f}(D)$ where $f=x(x+1)\in\operatorname{\mathbb{F}}_{3}[x]$ and $D=\{x,x+1\}$}. We have $S_{D}=\{x,2x,x+1,2(x+1)\}$. The graph $G_{f}(D)$ is a regular graph of degree $4.$
Figure 1. The gcd-graph $G_{x(x+1)}(\{x,x+1\})$
Outline
In this article we introduce and study some foundational properties of gcd-graphs defined over $\mathbb{F}_{q}[x]$. In Section 2, we recall some standard definitions in graph theory that will be used throughout the article. Section 3 investigates the question of when a gcd-graph is connected and anti-connected. (Recall that a graph is anti-connected if its complement is connected.) In Section 4, we provide the necessary and sufficient conditions for a gcd-graph to be bipartite. We note that our investigation provides additional insights into the properties of Cayley graphs. For example, in recent works [29, 30], the authors examine the connectedness and bipartiteness of a related class of graphs known as generalized Paley graphs. Section 5 explores another graph-theoretic property of gcd-graphs: their primeness property. Here, we describe the necessary and sufficient conditions for a gcd-graph to be prime. We then apply this result to give an explicit criterion for a gcd-graph to be prime under some mild assumptions on the set $D$. Section 6 examines the spectra of gcd-graphs using symmetric algebras and Ramanujan sums. We find it quite remarkable that the explicit formulas for these spectra are identical to those that appeared in the number field case. In Section 7, we provide some sufficient conditions for gcd-graphs to be perfect. Finally, in Section 8, we investigate whether a given graph can be realized as an induced subgraph of a gcd-graph. As a by-product, we prove an analogous result to a theorem of Erdős and Evans in the number field case.
Code
@@ 2. Background from graph theory
The complete graph $K_{n}$ is the graph on $n$ vertices which are pairwisely adjacent.
We now review two types of graph products that appear in our work, starting with the tensor product.
@@
Next, we recall the definition of the wreath product of graphs.
Definition 2.3 (Wreath product of graphs).
Let $G,H$ be two graphs. We define the wreath product of $G$ and $H$ as the graph GHG*H with the following data:
@@
    The vertex set of GHG*H is the Cartesian product $V(G)\times V(H)$,
@@
    $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent in GHG*H if either $(g,g^{\prime})\in E(G)$ or $g=g^{\prime}$ and $(h,h^{\prime})\in E(H)$.
Definition 2.4 (Graph morphisms).
Let $G_{1}$ and $G_{2}$ be two graphs. We define a graph morphism between $G_{1}$ and $G_{2}$ to be a map $f$ from $V(G_{1})$ to $V(G_{2})$ that preserves edges. More precisely, if $u,v\in V(G_{1})$ are adjacent in $G_{1}$, then $f(u),f(v)$ are adjacent in $G_{2}$.
Finally, we recall the definition of induced subgraphs.
Definition 2.5 (Induced subgraphs).
Let $G_{1}$ and $G_{2}$ be two graphs. We say that $G_{1}$ is an induced subgraph of $G_{2}$ if there exists a graph morphism $f\colon G_{1}\to G_{2}$ such that the following conditions hold.
@@
    For $\in G_{1}, u,vu,v$ are adjacent if and only if $f(u)$ and $f(v)$ are adjacent in $G_{2}.$
3. Connectedness of gcd-graphs
In this section, we will give the necessary and sufficient conditions for $G_{f}(D)$ to be connected. Unlike the number field case where these conditions are rather straightforward, the function field case is a bit more complicated. This is partially due to the fact that the unitary Cayley graph on $\operatorname{\mathbb{F}}_{q}[x]/(f)$ is not always connected. We start our discussion with the following simple observation.
@@ 3. Connectedness of gcd-graphs
Let $f_{0}=\gcd(f_{1},\ldots,f_{k})$. We observe that if two vertices $u$ and $v$ are adjacent then $f_{0}$ divides uvu-v. Now since $G_{f}(D)$ is connected, there is a path connecting 0 and $1$. This implies that $f_{0}$ divides 10=11-0=1. Hence $\gcd(f_{1},\ldots,f_{k})=1$. ∎
[…]
@@
Let $R=\operatorname{\mathbb{F}}_{q}[x]/(f)$. We need to show that $R=\langle S_{D}\rangle$ the abelian group generated by $S_{D}$. Let $a\in R.$ Because $\gcd(f_{1},f_{2},\ldots,f_{k})=1$, we can find $a_{1},a_{2},\ldots,a_{k}\in R$ such that
@@
Since $G_{f}(\{1\})$ is connected, for each $1\leq i\leq k$, we can find write
@@
Since $f_{i}\mid f$, we see that $\gcd(s_{ij}f_{i},f)=f_{i}\in D$. This shows that $a\in\langle S_{D}\rangle$. Since this is true for all $a$, we conclude that $R=\langle S_{D}\rangle$. ∎
@@
Proposition 3.3 (See [9, Lemma 4.33]).
$G_{f}(\{1\})$ is not connected if and only if $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$
@@
Let $g$ be a divisor of $f$ and $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}[x]/g$ be the canonical projection map. Let
$\overline{D}=\{\gcd(f_{i},g):1\leq i\leq k\}.$
@@
Let $\bar{a},\bar{b}$ be two vertices in $G_{g}(\overline{D}).$ Since $\Phi_{f,g}$ is surjective, we can find $a,b\in\operatorname{\mathbb{F}}_{q}[x]/(f)$} such that $\Phi_{f,g}(a)=\bar{a},\Phi_{f,g}(b)=\bar{b}.$ Since $G_{f}(D)$ is connected, there is a path $P$ from $a$ to $b.$ By 3.6, $\Phi_{f,g}(P)$ is a path from $\bar{a}$ to $\bar{b}.$ This shows that $G_{g}(\overline{D})$ is connected as well. ∎
@@
Suppose that $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$ Let
$\Phi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}[x]/(x(x+1))$
be the canonical projection map and $\bar{D}$ as described above. Then $G_{f}(D)$ is connected if and only if
@@
    the graph $G_{x(x+1)}(\bar{D})$ is connected. This condition is equivalent to D ∖{x(x+1) }2.
@@
By 3.7, we know that if $G_{f}(D)$ is connected then $(1)$ and $(2)$ holds. Conversely, let us assume that $(1)$ and $(2)$ both hold. We will show that $G_{f}(D)$ is connected. This is equivalent to showing that $R:=\operatorname{\mathbb{F}}_{q}[x]/(f)=\langle S_{D}\rangle.$
@@
We first claim that if $g\in R$ such that $x(x+1)\mid g$ then $g=s_{1}+s_{2}$ where $s_{1},s_{2}\in R^{\times}.$ For an element $\in R, ss$ is a unit if and only $\Phi^{\text{ss}}(s)\in(R^{\text{ss}})^{\times}$ where $R^{\text{ss}}=R/J(R)$ is the semisimplification of $R$ (see [9, Proposition 4.30]). Therefore, for this statement, we can assume that $R=R^{\text{ss}}$; namely $f$ is a squarefree polynomial. If we write $f=x(x+1)h$ where $\gcd(x(x+1),h)=1$ then we have the isomorphism
$R\cong\operatorname{\mathbb{F}}_{2}[x]/(x(x+1))\times\operatorname{\mathbb{F}}_{2}[x]/(h).$
Under this isomorphism, $g$ is sent to $(0,g_{1})$ where $g_{1}\in\operatorname{\mathbb{F}}_{2}[x]/(h).$ Since $F_{2}[x]/(h)$ is a product of fields of order bigger than $2$, every element in it can be written as the sum of two units; say $g_{1}=t_{1}+t_{2}$ where $t_{1},t_{2}\in(\operatorname{\mathbb{F}}_{2}[x]/(f))^{\times}.$ We then see that $g=s_{1}+s_{2}$ where $s_{1}=(1,t_{1})$ and $s_{2}=(1,t_{2})$. By definition $s_{1},s_{2}\in R^{\times}.$
@@
Since $s_{1i},s_{2i}\in R^{\times}$, $s_{1i}f_{i},s_{2i}f_{i}\in S_{D}.$ This shows that $g\in\langle S_{D}\rangle.$
Finally, let $g$ now be an arbitrary element in $R$. We claim that $g\in\langle S_{D}\rangle.$ By our assumption, the graph $G_{x(x+1)}(S_{\bar{D}})$ is connected, 0 and $\Phi_{f,x(x+1)}(g)$ are connected by a path. Consequently, we can write
@@
The proof of 3.8 also implies that for the case $\operatorname{\mathbb{F}}_{2}[x]$ and $x(x+1)\mid f$}, the map $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{2}[x]/(f)\to\operatorname{\mathbb{F}}_{2}[x]/(x(x+1))$ has the property that $a$ and $b$ belong to the same connected component in $G_{f}(D)$ if and only $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ belong to the same connected component in $G_{x(x+1)}(\overline{D}).$ In fact, by 3.7, if $a$ and $b$ are connected by a path then $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ are connected by a path. Conversely, if $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ are connected by a path in $G_{x(x+1)}(\overline{D})$, we can write
@@
where $n_{i}\in\operatorname{\mathbb{Z}}$ and $\gcd(h_{1},f)\in D$. From the proof of 3.8, we know that $x(x+1)R\subseteq\langle S_{D}\rangle$. Therefore, we conclude that $a-b\in\langle S_{D}\rangle.$ By definition, $a$ and $b$ are connected by a path in $G_{f}(D)$.
We conclude that $G_{f}(D)$ and $G_{x(x+1)}(\overline{D})$ has the same number of connected component. In particular, the number of connected components in $G_{f}(D)$ is at most $2.$
Example 3.10.
Fig. 2 shows the graph $G_{f}(D)$ where $f=x^{2}(x+1)\in\operatorname{\mathbb{F}}_{2}[x]$ and $D=\{x,x+1\}$}. We have $S_{D}=\{x,x+1,x^{2}+1\}$. The graph $G_{f}(D)$ is a regular graph of degree $3.$ Furthermore, by 3.8, it is connected.
Figure 2. The gcd-Cayley graph $G_{x^{2}(x+1)}(\{x,x+1\})$